# The inverse Kalman filter

### v3

@@ 4.4 Estimating cell-cell interactions on anisotropic substrates
where $n_{p}(\tau)$ is the cell count at time $\tau$, $l=1,2$ correspond to horizontal and vertical directions, respectively, and $\epsilon_{i^{\prime},l}(\tau)\sim\mathcal{N}(0,\sigma^{2}_{0,l}(\tau))$ denotes the noise. Inspired by the modified Vicsek model, we set $d_{k,l}=v_{k,l}(\tau-1)$. To account for velocity decay caused by increasing cell confluence, we model the noise variances as $\sigma^{2}_{0,l}(\tau)=\omega_{l}\sigma^{2}_{v,l}(\tau-1)$, where $\sigma^{2}_{v,l}(\tau-1)$ is the sample velocity variance at time $(\tau-1)$, and $\omega_{l}$ is a parameter estimated from data for $l=1,2$. The neighboring set $ne_{i^{\prime},l}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r_{l}\mbox{ and }\mathbf{v}_{i^{\prime}}(\tau-1)\cdot\mathbf{v}_{k}(\tau-1)>0\}$, where $r_{l}$ denotes the interaction radius of direction $l$ for $l=1,2$, excludes particles moving in opposite directions, reflecting the observed gliding and intercalating behavior of cells [36].
Figure 8: The blue dashed curves show the predicted interaction function for the horizontal and vertical directions. The grey shaded area is the 95% posterior credible interval. The purple line of slope 1 is the prediction using the Vicsek model. The light pink histogram shows the velocity distribution of all training samples.
@@
Our nonparametric estimation of the interaction functions is compared with two models for one-step-ahead forecasts of directional velocities: the original Vicsek model introduced in Section 4.2 and the anisotropic Vicsek model, which predicts velocities using the average velocity of neighboring cells from the previous time frame. As shown in Table 2, our model outperforms both Vicsek models in RMSE for one-step-ahead forecasts of directional velocities, despite the large inherent stochasticity in cellular motion. Moreover, our model has notably shorter average interval lengths that cover approximately $95\%$ of the held-out observations. These findings underscore the importance of the IFK-CG algorithm in enabling the use of large datasets to overcome high stochasticity and capture the underlying dynamics of alignment processes.
@@
Baseline Vicsek $4.5\times 10^{-3}$ 0.011 80% $5.3\times 10^{-3}$ 0.011 95%
Anisotropic Vicsek $3.9\times 10^{-3}$ 0.024 98% $2.6\times 10^{-3}$ 0.020 99%
@@ 5 Concluding remarks
Several future directions are worth exploring. First, computing large matrix-vector products is ubiquitous, and the approach can be extended to different models, including some large neural network models. Second, it is an open topic to scalably compute the logarithm of the determinant of the covariance in Eq. (3) without approximation. Third, the new approach can be extended to estimate latent functions when forward equations are unavailable or computationally intensive in nonlinear or non-Gaussian dynamical systems, where ensemble Kalman filter [6, 33] and particle filter [20] are commonly used. Fourth, the proposed approach can be extended for estimating and predicting multivariate time series [21] and generalized factor processes for categorical observations [4].
@@ Acknowledgement
Xinyi Fang acknowledges partial support from the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487. We thank the editor, associate editor, and three anonymous referees for their comments that substantially improved this article.
@@
The Supplementary Material contains (i) proofs of all lemmas; (ii) a summary of the connection between GPs with Matérn covariance (roughness parameters 0.5 and 2.5) and DLMs, with closed-form $\mathbf{F}_{t,j}$, $\mathbf{G}_{t,j}$ and $\mathbf{W}_{t,j}$; (iii) the conjugate gradient algorithm; (iv) procedures for the scalable computation of the predictive distribution in particle dynamics; (v) an application of the IKF-CG algorithm for predicting missing values in lattice data; (vi) the parameter estimation method; (vii) computational complexity analysis; (viii) additional numerical results for particle interaction; and (ix) numerical results for predicting missing values in lattice data.
@@ Appendix A. Kalman Filter
The Kalman filter [16] for dynamic linear models in (1)-(2) is summarized in Lemma 5.
Lemma 5.

### v4

@@ 4.4 Estimating cell-cell interactions on anisotropic substrates
where $n_{p}(\tau)$ is the cell count at time $\tau$, $l=1,2$ correspond to horizontal and vertical directions, respectively, and $\epsilon_{i^{\prime},l}(\tau)\sim\mathcal{N}(0,\sigma^{2}_{0,l}(\tau))$ denotes the noise. Inspired by the modified Vicsek model, we set $d_{k,l}=v_{k,l}(\tau-1)$. To account for velocity decay caused by increasing cell confluence, we model the noise variances as $\sigma^{2}_{0,l}(\tau)=\omega_{l}\sigma^{2}_{v,l}(\tau-1)$, where $\sigma^{2}_{v,l}(\tau-1)$ is the sample velocity variance at time $(\tau-1)$, and $\omega_{l}$ is a parameter estimated from data for $l=1,2$. The neighboring set $ne_{i^{\prime},l}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r_{l}\mbox{ and }\mathbf{v}_{i^{\prime}}(\tau-1)^{T}\mathbf{v}_{k}(\tau-1)>0\}$, where $r_{l}$ denotes the interaction radius of direction $l$ for $l=1,2$, excludes particles moving in opposite directions, reflecting the observed gliding and intercalating behavior of cells [37].
Figure 8: The blue dashed curves show the predicted interaction function for the horizontal and vertical directions. The grey shaded area is the 95% posterior credible interval. The purple line of slope 1 is the prediction using the unnormalized Vicsek model. The light pink histogram shows the velocity distribution of all training samples.
@@
Our nonparametric estimation of the interaction functions is compared with two models for one-step-ahead forecasts of directional velocities: the unnormalized Vicsek model introduced in Section 4.2, which uses a shared interaction radius for both velocity components, and the anisotropic Vicsek model, which uses distinct radii in each direction. As shown in Table 2, both Vicsek models have similar RMSE as the difference of interaction radii does not have a large impact for estimating interaction functions. Our model outperforms both Vicsek models in RMSE for one-step-ahead forecasts of directional velocities, despite the large inherent stochasticity in cellular motion. Moreover, our model has notably shorter average interval lengths that cover approximately $95\%$ of the held-out observations. These findings underscore the importance of the IFK-CG algorithm in enabling the use of large datasets to overcome high stochasticity and capture the underlying dynamics of cell alignment processes.
@@
Unnormalized Vicsek $3.9\times 10^{-3}$ 0.029 99% $2.6\times 10^{-3}$ 0.029 99%
Anisotropic Vicsek $3.9\times 10^{-3}$ 0.034 99% $2.6\times 10^{-3}$ 0.022 99%
@@ 5 Concluding remarks
Several future directions are worth exploring. First, computing large matrix-vector products is ubiquitous, and the approach can be extended to different models, including some large neural network models. Second, it is an open topic to scalably compute the logarithm of the determinant of the covariance in Eq. (3) without approximation. Third, the new approach can be extended to estimate latent functions when forward equations are unavailable or computationally intensive in nonlinear or non-Gaussian dynamical systems, where ensemble Kalman filter [6, 34] and particle filter [21] are commonly used. Fourth, our approach can be extended for estimating and predicting multivariate time series [22] and generalized factor processes for categorical observations [4]. Finally, we aim to generalize the fast approach to scenarios with input dimensions larger than one.
@@ Acknowledgement
Xinyi Fang acknowledges partial support from the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487. Mengyang Gu acknowledges partial support from the National Science Foundation under Award No. OAC-2411043 and the University of California Multicampus Research Programs and Initiatives under Award No. M23PL599. We thank the editor, associate editor, and three anonymous referees for their comments that substantially improved this article.
@@
The Supplementary Material contains (i) proofs of all lemmas; (ii) a summary of the connection between Gaussian processes with Matérn covariance (roughness parameters 0.5 and 2.5) and DLMs, with closed-form expressions of $\mathbf{F}_{t,j}$, $\mathbf{G}_{t,j}$ and $\mathbf{W}_{t,j}$; (iii) the conjugate gradient algorithm; (iv) procedures for the scalable computation of the predictive distribution in particle dynamics; (v) an application of the IKF-CG algorithm for predicting missing values in lattice data; (vi) parameter estimation methods; (vii) computational complexity analysis; (viii) additional simulation results for estimating particle interaction; and (ix) numerical results for predicting missing values in lattice data.
@@ Appendix A. Kalman Filter
The Kalman filter [17] for dynamic linear models in (1)-(2) is summarized in Lemma A1.
Lemma A1.